OpenStudy (anonymous):

prove the identity (1+tan2x)/ (sin2x+cos2x)=sec2x

myininaya (myininaya):

is it (tanx)^2 or tan(2x)

myininaya (myininaya):

ok i think its tan(2x)

myininaya (myininaya):

ok tan(2x)=sin(2x)/cos(2x)

OpenStudy (anonymous):

tan2 x

OpenStudy (anonymous):

This question came up already, check for spanishkb31

OpenStudy (anonymous):

the identity works if the 2s are ^2

myininaya (myininaya):

multiply both top and bottom by cos(2x)

OpenStudy (anonymous):

sin^2(x) + cos^2(x) = 1

myininaya (myininaya):

so we have [cos(2x)+sin(2x)]/[cos(2x){sin(2x)+cos(2x)}]=1/cos(2x)=sec(2x)

OpenStudy (anonymous):

...leaving 1+tan^2(x) = sec^2(x) which is an identity, but can be further proved by multiplying through by cos^2(x): sin^2(x) + cos^2(x) = 1

OpenStudy (anonymous):

dear god thank you

myininaya (myininaya):

it doesnt have to be the "squareys"

OpenStudy (anonymous):

Haha this is a coincidence. It works in both cases; if 2 is the exponent or part of the angle.

myininaya (myininaya):

right :)